OpenStudy (anonymous):

estimate 43 7/8 divided bye 10 1/3

OpenStudy (anonymous):

hey, you spelled estimate correctly :D

OpenStudy (anonymous):

haha hilariouse -_-

OpenStudy (anonymous):

:{ lol, i'm doing the problem atm want to see work, or do u not care?

OpenStudy (anonymous):

i dont care lol ((:

OpenStudy (anonymous):

351/8 * 3/31 1053/248 4 61/248

OpenStudy (anonymous):

31/3

OpenStudy (anonymous):

you realize you just put 1053/248/24? you need an LCD

OpenStudy (anonymous):

gaara learn to count then come what is 43/10?

OpenStudy (anonymous):

It's about 44/10 is about 4.4

OpenStudy (anonymous):

learn that 10 1/3 does NOT make 3/31

OpenStudy (anonymous):

yeah it's 4 61/248 your work just got some typos

OpenStudy (anonymous):

learn that (a/b)/(c/d) = a/b * d/c

OpenStudy (anonymous):

Alternatively it's 43.875 /10.333.....

OpenStudy (anonymous):

lol i love this random estudier guy xD

OpenStudy (anonymous):

Depends what u mean by "estimate"

OpenStudy (anonymous):

gaara do you realize that 31/3 = 10 1/3?

OpenStudy (anonymous):

yes, i already put that if you would scroll up when you put your work down, you just left out a step

OpenStudy (anonymous):

lol i didnt say that that was the answer i was showing you your work

OpenStudy (anonymous):

@rainbowink do you have choices for this problem or no

OpenStudy (anonymous):

"351/8 * 3/31 1053/248 4 61/248" i'm tlaking about this

OpenStudy (anonymous):

tlaking?

OpenStudy (anonymous):

yes, i'm sure you can decipher a typo.

OpenStudy (anonymous):

ok gaara can you show me how to do it step by step please

OpenStudy (anonymous):

*face palm* maybe if you weren't so conceded you would have read that i stated the steps were right, BUT you left out one.

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

351/8 divided by 31/3 you just jumped into the multiplication, which is what i don't understand you are putting down the steps, yet you skip around